OpenStudy (yanni):

y=-6/19x+13/4 find the y-intercept

OpenStudy (australopithecus):

sorry I mean set x = 0 and solve for y

OpenStudy (saifoo.khan):

Whenever y-intercepts, x = 0

OpenStudy (saifoo.khan):

\[y = -\frac{6}{19} (0) + \frac{13}{4}\]

OpenStudy (australopithecus):

so y = 6/19(0) + 13/4 you cannot divide by zero so you need to rearrange the equation

OpenStudy (australopithecus):

thus there are no y intercepts

OpenStudy (australopithecus):

You can tell by looking at the graph https://www.wolframalpha.com/input/?i=y%3D-6%2F19x%2B13%2F4